# Universality in many-body driven systems with an umbilic point

### v1

@@ Abstract
I Introduction
One-dimensional driven stochastic systems of particles interacting with a short range interactions with one global conservation law belong to the renowned Kardar-Parisi-Zhang (KPZ) universality class [1], with the exact scaling function found by Prähofer and Spohn [2]. KPZ universality is expected under rather general assumptions which include a finite steady-state correlation length, stationarity of spatially homogeneous distributions, a presence of noise and of long-lived modes stemming from the conservation law. Physically, realizations of KPZ universality have been demonstrated in a series of beautiful experiments by Takeuchi [7]. Dynamical behavior of many-body systems with the two or more long-lived modes at large space and time scales hase been shown to be well described by nonlinear fluctuating hydrodynamics and mode-coupling theory [3]. A key assumption of the resulting nonlinear fluctuating hydrodynamics and mode coupling analysis is a spatial separation of long-lived modes with time, meaning that velocities of the modes are all different. Such an assumption of strict hyperbolicity is also a cornerstone of the Lax theory of shocks [4].
@@ I Introduction
A recent paper [13] treats the question of universality of a weakly hyperbolic system with two conserved quantities using two approaches: firstly, by numerical integration of a system of two stochastic Burgers equations, and secondly, by Monte Carlo simulation of a two-lane bidirectional particle model with an umbilic point. In particular, the scaling of space-time correlations in those two systems is investigated to infer the dynamical exponent and the distribution of the time-integrated currents. An important feature of both models is so-called cyclicity condition for the mode-coupling matrices $G^{\alpha}$ which guarantees time-stationarity of the Gaussian measure. For both cases the authors find a match between the space-time correlators for both lattice and continuous model. The authors of [13] conclude that the dynamical exponents for the umbilic modes are KPZ-like, $z_{1}=z_{2}=3/2$, as in the the single-component Burgers equation, and that the scaling functions are likely to be coupling-dependent.
@@
The purpose of the present communication is to present evidence that not only the scaling functions but also the dynamical exponents for the umbilic modes are distinct from the KPZ scaling exponents, $z_{1},z_{2}\neq 3/2$. This contradicts one of main statements of [13]. Namely, our results strongly suggest that both scaling functions and dynamical scaling exponents are coupling-dependent. We show that the reason for the contradiction is the short time window employed in [13], based on which the observations and the conclusions were made. We have performed Monte-Carlo simulations over much larger time scale $t\leq 10^{5},L=5\times 10^{5}$ for the two-lane lattice model, compared to $160008000<t<$ 16000, $L=2^{16}\approx 6.5\times 10^{4}$ in [13]. In view of the above, umbilic universality appears to be an even more exciting and challenging topic which needs further studies. Indeed, our findings are in sharp contrast with the strictly hyperbolic case where a discrete family of the dynamical exponents (Fibonacci universality classes [10]) exhaust all possible values of the dynamical exponents of a model with two conserved quantities.
II Description of an umbilic point. Saddle and convex topologies
@@
We consider a two-lane bidirectional model of left-moving and right-moving particles, with hardcore interaction between them (each site cannot be occupied by more than one particle) proposed in [12], a schematic representation of which is given in Fig. 1. The exchange rules (hopping rates) of the left movers are obtained from those of the right movers by a reflection. The model has one inter-chain interaction parameter $\gamma\in[0,\infty)$. For $\gamma=1$ the system decouples into two separate totally asymmetric exclusion processes, characterized by KPZ universality.
@@
The invariant measure of the bidirectional model factorizes over pairs of vertical sites, thus allowing to obtain an explicit expression for steady microscopic currents $j_{1},j_{2}$ as a function of the average particle densities on upper and lower lanes $0\leq u,v\leq 1$. The eigenvalues of the flux Jacobian,
${\cal D}_{J}(u,v)$ $=\left(\begin{array}[]{cc}\frac{\partial j_{1}}{\partial u}&\frac{\partial j_{1}}{\partial v}\\ \frac{\partial j_{2}}{\partial u}&\frac{\partial j_{2}}{\partial v}\end{array}\right),$ (3)
${\cal D}_{J}(u,v)\psi_{\alpha}(u,v)$ $=c_{\alpha}(u,v)\psi_{\alpha}(u,v),$ (4)
[…]
@@
In case of non-zero characteristic umbilic point velocity $c_{1}(u_{0},v_{0})=c_{2}(u_{0},v_{0})=c$, the same reasoning applies after a renormalization $c_{\alpha}(u,v)\rightarrow c_{\alpha}(u,v)-c$. In other words, an umbilic point at $u_{0},v_{0}$ with $c_{1}(u_{0},v_{0})=c_{2}(u_{0},v_{0})=c$ is isolated if there exists an $\epsilon_{0}>0$ such that for any $0<\epsilon<\epsilon_{0}$, and $0\leq\varphi<2\pi$, $c_{\alpha}(u_{0}+\epsilon\cos\varphi,v_{0}+\epsilon\sin\varphi)\neq c$, $\alpha=1,2$.
@@
In a system with $K>2$ conservation laws further topological possibilities appear, since the phase space becomes $K$-dimensional.
@@
Ref. [13] shows that the bidirectional model Fig. 1 at the unique umbilic point $u=v=1/2$, for any coupling $\gamma$ corresponds, in the framework of mode coupling theory, to a system of coupled stochastic Burgers equations with cyclic coupling matrices and a degenerate flux Jacobian. Moreover, for the critical value of the coupling $\gamma_{\rm crit}=1/4$ (the $\gamma$ value separating the isolated and non-isolated umbilic point scenario) the mode coupling matrices become especially simple, and correspond to a dynamics, maximally distant from the decoupled scalar KPZ growth processes.
III Numerical investigation of the dynamical structure factor
@@
We now investigate the space-time correlation matrix (dynamical structure factor) of the process. To make the connection with a system of coupled stochastic Burgers equations we choose a basis where the static covariance matrix becomes an identity matrix [13], see also Appendix A. In this basis the space-time correlation matrix $S_{\alpha\beta}$ becomes also diagonal, with the diagonal elements given by
$S_{11}(x,t)$ $=(1+\gamma^{1/2})\cdot$ (5)
$\left(\tilde{S}_{11}(x,t)+\tilde{S}_{22}(x,t)+\tilde{S}_{12}(x,t)+\tilde{S}_{21}(x,t)\right)$
$S_{22}(x,t)$ $=(1+\gamma^{-1/2})\cdot$ (6)
$\left(\tilde{S}_{11}(x,t)+\tilde{S}_{22}(x,t)-\tilde{S}_{12}(x,t)-\tilde{S}_{21}(x,t)\right)$
where
$\tilde{S}_{\alpha\alpha}(x=ka,t)=\langle\eta_{k}^{\alpha}(t)\eta_{0}^{\alpha}(0)\rangle,$ (7)
@@
are “bare” space-time correlation matrices of the two-lane Markov process (with particle densities $\eta_{x}^{\alpha}(t)$) in Fig. 1 where $a$ is the lattice spacing and $k$ the lattice position. Up to a rescaling, the functions $S_{\alpha\alpha}(x,t)$ are identical to those studied in [13]. According to the scaling hypothesis, at large space and time scales one expects
$S_{\alpha\alpha}(x,t)\sim t^{-1/z_{\alpha}}f_{\alpha}\left(xt^{-1/z_{\alpha}}\right),$ (8)
where $z_{\alpha}$ is a dynamical exponent.
@@
To quantify the scaling (8) we have performed extensive Monte Carlo simulations for a periodic system of size $L=5\times 10^{5}$ up to $t=10^{5}$ time units, starting from spatially homogeneous states with densities $1/2$ on both lanes. Measurement were done at all sites in parallel, since due to translational invariance $S_{\alpha\alpha}(x,t)$ depends only on the distance between two points. Further details on numerical simulations can be found in the Appendix B. Fig. 3 shows the main quantities of interest, the estimated dynamical exponents $z_{1}(t),z_{2}(t)$ as functions of time; according to the scaling hypothesis they should converge to stationary values, $\lim_{t\to\infty}z_{\alpha}(t)=z_{\alpha}$.
Both the data collapse of $S_{\alpha\alpha}(x,t)$ at different times, see Fig. 5, and the late time behavior of $z_{\alpha}(t)$ suggest the validity of the scaling (8), with $z_{\alpha}$
@@
$z_{2}=1.587\pm 0.005$ (9)
where the $z_{\alpha}$ are $0.5\%$ accurate. Most notably, both $z_{1}$ and $z_{2}$ are distinct from the KPZ exponent $z_{\rm KPZ}=3/2$, suggested in [13]. To emphasize our point, in Fig.  4 we plot our data for $S_{11},S_{22}$ in a form which should show data collapse if the hypothesis $z_{1}=3/2$ was valid. Clearly, the data collapse in Fig. 4 is poor, invalidating the hypothesis. In contrast, Fig. 5 shows asymptotic data collapse for the values $z_{1},z_{2}$ in (9).
Figure 3: Lower Panel: $z_{1}(t)$ and $z_{2}(t)$. Upper Panel: scaling of the maxima $\max S_{\alpha\alpha}(t)$ of the distribution: $\max S_{\alpha\alpha}(t)\sim t^{-1/z}$ according to scaling hypothesis. Parameters: $L=5\times 10^{5}$, $\tau=0$, $M=1$, $P=2500$, $R=50$, and $p=0.01$ see Appendix B for details.
Figure 4: Rescaled $S_{11},S_{22}$ versus rescaled space, for different times. If the hypothesis $z_{1}=z_{2}=3/2$ was valid the above data should show data collapse. System parameters: $L=2.5\times 10^{5}$, $\tau=80$, $M=10^{4}$, $P=983$, $R=1$, see Appendix B for details.
Figure 5: Rescaled $S_{11},S_{22}$ versus rescaled space, for different times. The best data collapse for large times is achieved with $z_{1}=1.456$, $z_{2}=1.587$. Parameters: $L=2.5\times 10^{5}$, $\tau=80$, $M=10^{4}$, $P=983$, $R=1$, see Appendix B for details
III.1 Other values of the coupling
To support our findings, we have investigated the space-time correlator $S_{\alpha\alpha}(x,t)$ at different values of the coupling parameter $\gamma$. Only for $\gamma=1$, the point where the dynamics splits into two independent totally asymmetric processes, do we retrieve the well-known KPZ scaling of a stochastic scalar Burgers equation $z_{1}=z_{2}\equiv 3/2$, with $f_{\alpha}=f_{\rm KPZ}$ the celebrated KPZ scaling function obtained in Ref. [2]. For all other value of the coupling parameter $\gamma<1$ we find not only scaling functions distinct from $f_{\rm KPZ}$ but also different dynamical exponents, see Fig. 6. We observe a clearly tendency of both $z_{1},z_{2}\rightarrow 3/2$ as $\gamma$ increases towards $\gamma=1$, the point of decoupling into separate simple exclusion processes. Also the shape of scaling functions $f_{1},f_{2}$ become closer to KPZ as $\gamma$ increases towards $1$, data not shown. The data for the regime $\gamma<\gamma_{\rm crit}$ which correspond to an isolated umbilic point ( saddle point topology) are less transparent. The critical value of the coupling $\gamma=1/4$, at which the saddle point appears, corresponds to the maximal $|z_{1}-z_{2}|$ difference. The limit $\gamma\rightarrow 0$ cannot be reached via Monte Carlo simulations since the effective MC dynamics close to $\gamma=0$ at the umbilic point becomes too slow. Further studies of this regime are warranted.
@@ III.1 Other values of the coupling
IV Conclusions
We have presented a numerical study of the dynamical structure factor for a many-body particle system with an umbilic point and two conservation laws. Both isolated and non-isolated umbilic point were considered. We have investigated the validity of the scaling hypothesis (8) in a basis which diagonalizes both the dynamical structure matrix $S_{\alpha\beta}(x,t)$ and the static covariance matrix. On the basis of the data we conclude that, within the hypothesis (8) that: (1) the dynamical exponents $z_{1},z_{2}$ and (2) the dynamical scaling functions $f_{1},f_{2}$ vary continuously with the parameter of the model $\gamma$ in a nontrivial way, i.e. the scaling functions are not merely dilated. The second point is in agreement with an earlier study [13].

### v2

@@ Abstract
pacs:
1 Introduction
One-dimensional driven stochastic systems of particles interacting with a short range interactions with one global conservation law belong to the renowned Kardar-Parisi-Zhang (KPZ) universality class [1], with the exact scaling function found by Prähofer and Spohn [2]. KPZ universality is expected under rather general assumptions which include a finite steady-state correlation length, stationarity of spatially homogeneous distributions, a presence of noise and of long-lived modes stemming from the conservation law. Physically, realizations of KPZ universality have been demonstrated in a series of beautiful experiments by Takeuchi et al [3, 4]. Dynamical behavior of many-body systems with the two or more long-lived modes at large space and time scales hase been shown to be well described by nonlinear fluctuating hydrodynamics and mode-coupling theory [5]. A key assumption of the resulting nonlinear fluctuating hydrodynamics and mode coupling analysis is a spatial separation of long-lived modes with time, meaning that velocities of the modes are all different. Such an assumption of strict hyperbolicity is also a cornerstone of the Lax theory of shocks [6].
@@ 1 Introduction
A recent paper [10] treats the question of universality of a weakly hyperbolic system with two conserved quantities using two approaches: firstly, by numerical integration of a system of two stochastic Burgers equations, and secondly, by Monte Carlo simulation of a two-lane bidirectional particle model with an umbilic point. In particular, the scaling of space-time correlations in those two systems is investigated to infer the dynamical exponent and the distribution of the time-integrated currents. An important feature of both models is a so-called cyclicity condition for the mode-coupling matrices $G^{\alpha}$ which guarantees time-stationarity of the Bernoulli measure. For both cases the authors find a match between the space-time correlators for both lattice and continuous model. The authors of [10], relying on the available direct numerical simulations, conclude that the dynamical exponents for the umbilic modes are KPZ-like, $z_{1}=z_{2}=3/2$, as in the the single-component Burgers equation, and that the scaling functions are coupling-dependent.
@@
The purpose of the present communication is to present evidence that not only the scaling functions but also the dynamical exponents for the umbilic modes are distinct from the KPZ scaling exponents, $z_{1},z_{2}\neq 3/2$. This contradicts one of main statements of [10]. Namely, our results strongly suggest that both scaling functions and dynamical scaling exponents are coupling-dependent. We show that the reason for the contradiction is the short time window employed in [10], based on which the observations and the conclusions were made. We have performed Monte-Carlo simulations over much larger time scale $t\leq 10^{5},L=5\times 10^{5}$ for the two-lane lattice model, compared to $160008000<t<$ 16000, $L=2^{16}\approx 6.5\times 10^{4}$ in [10], see D for a comparison of the data. Our numerical simulation of the system of coupled Burgers equations studied in [10] similarly shows deviations from $z_{1},z_{2}\neq 3/2$ at longer times. In view of the above, umbilic universality appears to be an even more exciting and challenging topic which needs further studies. Indeed, our findings are in sharp contrast with the strictly hyperbolic case where a discrete family of the dynamical exponents (Fibonacci universality classes [7]) exhaust all possible values of the dynamical exponents of a model with two conserved quantities.
2 Description of an umbilic point. Saddle and convex topologies
@@
We consider a two-lane bidirectional model of left-moving and right-moving particles, with hardcore interaction between them (each site cannot be occupied by more than one particle) proposed in [11], a schematic representation of which is given in Fig. 1. A blue particle at site $k$ in the upper lane can hop to the right neighbouring site $k+1$ provided it is empty, with the rates shown in Fig. 1, which depend on the occupation of sites $k,k+1$ in the other (lower) lane. Likewise, on the lower lane (red particles) can hop to the left, with the hopping rates obtained from those of the right movers by a reflection. The model contains one inter-chain interaction parameter $\gamma\in[0,\infty)$, and is a reduction of a more general two-lane exclusion process, proposed in [11]. For $\gamma=1$ the system decouples into two separate totally asymmetric exclusion processes, characterized by KPZ universality.
The time evolution of the model is described by a continuous-time Markov process with above described rates. It was shown in [11] that the stationary state probabilities $P_{n_{1},n_{2},\ldots,n_{L}}^{m_{1},m_{2},\ldots,m_{L}}$ of configurations on a periodic lattice of length $L$ is given by
$P_{n_{1},n_{2},\ldots,n_{L}}^{m_{1},m_{2},\ldots,m_{L}}=Z^{-1}\prod_{k=1}^{L}e^{-\nu n_{k}m_{k}}$ (2.1)
@@
where $e^{\nu}=\gamma$ and $n_{k},m_{k}$ are occupation numbers $n_{k},m_{k}=0,1$ for the lower and upper chain respectively. The factorization (2.1) allows one to obtain explicit expressions for steady microscopic currents $j_{1},j_{2}$ as a function of the average particle densities on upper and lower lanes $0\leq u,v\leq 1$. The eigenvalues of the flux Jacobian,
${\cal D}_{J}(u,v)$ $=\left(\begin{array}[]{cc}\frac{\partial j_{1}}{\partial u}&\frac{\partial j_{1}}{\partial v}\\ \frac{\partial j_{2}}{\partial u}&\frac{\partial j_{2}}{\partial v}\end{array}\right),$ (2.4)
${\cal D}_{J}(u,v)\psi_{\alpha}(u,v)$ $=c_{\alpha}(u,v)\psi_{\alpha}(u,v),$ (2.5)
$c_{1}(u,v)$ and $c_{2}(u,v)$ are real continuous functions and correspond to characteristic velocities, i.e. velocities of local perturbations above stationary state with densities $u,v$. At $u=v=1/2$ the flux Jacobian becomes degenerate, $c_{1}=c_{2}=0$, yielding a unique umbilic point in the $u,v$ plane (all other $u,v$ points correspond to $c_{1}\neq c_{2}$). The phase space $0\leq u,v\leq 1$ can consequently be divided into regions the $G_{\pm\pm}$ and $G_{-+}$ according to signs of $c_{1}$ and $c_{2}$ respectively, e.g. $G_{-+}$ is the region where $c_{1}<0$ while $c_{2}>0$. The umbilic point $G_{00}$ at the center $u=v=1/2$ is either a point of contact of all three regions regions for $\gamma>\gamma_{crit}=1/4$ or an isolated point inside the $G_{-+}$ phase for $40<\gamma<1/$ 4 (in this case, $G_{++}$ and $G_{--}$ are separated, see Fig. 2). The two cases $\gamma>\gamma_{crit}$ and $\gamma<\gamma_{crit}$ thus correspond to distinct topologies. Indeed, the surface topology of the currents $j_{\alpha}(u,v)$ at the umbilic point is convex for $\gamma_{crit}\geq\gamma$ and forms a saddle for $\gamma<\gamma_{crit}$, see [8].
@@
In case of non-zero characteristic umbilic point velocity $c_{1}(u_{0},v_{0})=c_{2}(u_{0},v_{0})=c$, the same reasoning applies after a renormalization $c_{\alpha}(u,v)\rightarrow c_{\alpha}(u,v)-c$. In other words, an umbilic point at $u_{0},v_{0}$ with $c_{1}(u_{0},v_{0})=c_{2}(u_{0},v_{0})=c$ is isolated if there exists an $\epsilon_{0}>0$ such that for any $0<\epsilon<\epsilon_{0}$, and $0\leq\varphi<2\pi$, $c_{\alpha}(u_{0}+\epsilon\cos\varphi,v_{0}+\epsilon\sin\varphi)\neq c$, $\alpha=1,2$. In a system with $K>2$ conservation laws further topological possibilities appear, since the phase space becomes $K$-dimensional.
@@
Ref. [10] shows that the bidirectional model at the unique umbilic point $u=v=1/2$ and for any coupling $\gamma$ corresponds (in the framework of Taylor expansion to the second order of the stationary currents as function of particle densities) to a system of coupled stochastic Burgers equations with cyclic coupling matrices and a degenerate flux Jacobian. Moreover, for the critical value of the coupling $\gamma_{\rm crit}=1/4$ (the $\gamma$ value separating the isolated and non-isolated umbilic point scenario) the mode coupling matrices become especially simple, and correspond to a dynamics, maximally distant from a pair of decoupled scalar KPZ growth processes.
Figure 3: Rescaled bare diagonal $\tilde{S}_{11}$ (upper curves) and off-diagonal $\tilde{S}_{12}$ (lower curves) correlators of the two-lane process at different times rescalied with the dynamical exponents $z_{1}=z_{2}=3/2$. System parameters: $L=5\times 10^{5}$, $\tau=80$, $M=10^{4}$, $P=983$, $R=1$, see B for details.
3 Numerical investigation of the dynamical structure factor
We first investigate the “bare” space-time correlation matrix (dynamical structure factor) of the two-lane Markov process
$\tilde{S}_{\alpha\alpha}(x=ka,t)=\langle\eta_{k}^{\alpha}(t)\eta_{0}^{\alpha}(0)\rangle,$ (3.1)
@@ 3 Numerical investigation of the dynamical structure factor
with particle densities $\eta_{x}^{\alpha}(t)$ in Fig. 1 where $\alpha$ is the lane index, $a$ the lattice spacing and $k$ the lattice position. To facilitate the comparison to other models and establish a connection with the system of coupled stochastic Burgers equations of Ref. [10] we also consider a basis where the static covariance matrix becomes an identity matrix, see A. In this basis the space-time correlation matrix $S_{\alpha\beta}$ becomes diagonal, with the diagonal elements given by
$S_{11}(x,t)$ $=(1+\gamma^{1/2})\left(\tilde{S}_{11}(x,t)+\tilde{S}_{22}(x,t)+\tilde{S}_{12}(x,t)+\tilde{S}_{21}(x,t)\right)$ (3.2)
$S_{22}(x,t)$ $=(1+\gamma^{-1/2})\left(\tilde{S}_{11}(x,t)+\tilde{S}_{22}(x,t)-\tilde{S}_{12}(x,t)-\tilde{S}_{21}(x,t)\right)$ (3.3)
@@
Up to a rescaling, the functions $S_{\alpha\alpha}(x,t)$ are identical to those studied in [10]. According to the scaling hypothesis, at large space and time scales one expects
$S_{\alpha\alpha}(x,t)\sim t^{-1/z_{\alpha}}f_{\alpha}\left(xt^{-1/z_{\alpha}}\right),$ (3.4)
@@
where $z_{\alpha}$ is a dynamical exponent. To quantify the scaling (3.4) we have performed extensive Monte Carlo simulations for a periodic system of size $L=5\times 10^{5}$ up to $t=10^{5}$ time units, starting from spatially homogeneous states with densities $1/2$ on both lanes. Measurement were done at all sites in parallel, since $S_{\alpha\alpha}(x,t)$ depends only on the distance between two points by translational invariance. Further details on numerical simulations can be found in the B.
Figure 4: Upper panel: Scaling of the maxima $\max S_{\alpha\alpha}(t)$ of the distribution, $\max S_{\alpha\alpha}(t)\sim t^{-1/z_{\alpha}}$ according to scaling hypothesis. Lower panel: Dynamical exponents $z_{1}(t)$ and $z_{2}(t)$ as functions of time. Parameters: $L=5\times 10^{5}$, $\tau=0$, $M=1$, $P=2500$, $R=50$, and $p=0.01$ see B for details.
In chosen states the structure factor is invariant under exchange of lanes $\tilde{S}_{\alpha\beta}=\tilde{S}_{\beta\alpha}$, leaving two independent bare correlators shown in Fig. 3 rescaled with $z_{1}=z_{2}=3/2$. While the rescaled data for the diagonal correlator $\tilde{S}_{11}$ nearly collapses for different times, the off-diagonal correlator $\tilde{S}_{12}$ manifestly do not. We instead observe a gradual deepening of the central minimum accompanied by thickening tails due to the sum rule, precluding a uniform scaling collapse.
Figure 5: Left column: Rescaled $S_{11},S_{22}$ versus rescaled space, for different times. If the hypothesis $z_{1}=z_{2}=3/2$ was valid the data should show data collapse. Right column: The best data collapse for large times is achieved with $z_{1}=1.456$, $z_{2}=1.587$. System parameters as in Fig. 3.
We next consider the scaling of the correlators in the diagonal basis. Fig. 4 shows the main quantities of interest for quantifying the scaling (3.4), the estimated dynamical exponents $z_{1}(t),z_{2}(t)$ as functions of time; according to the scaling hypothesis they should converge to stationary values, $\lim_{t\to\infty}z_{\alpha}(t)=z_{\alpha}$. Both the data collapse of $S_{\alpha\alpha}(x,t)$ at different times, see right column of Fig. 5, and the late time behavior of $z_{\alpha}(t)$ suggest the validity of the scaling (3.4), with $z_{\alpha}$
@@
$z_{2}=1.587\pm 0.005$ (3.5)
where the $z_{\alpha}$ are $0.5\%$ accurate. Most notably, both $z_{1}$ and $z_{2}$ are distinct from the KPZ exponent $z_{\rm KPZ}=3/2$, suggested in [10]. To emphasize our point, in the left column of Fig. 5 we also plot our data for $S_{11},S_{22}$ in a form which should show data collapse if the hypothesis $z_{1}=z_{2}=3/2$ was valid. Clearly, the data collapse in the left column of Fig. 5 is poor, invalidating the hypothesis. By contrast, the right column of Fig. 5 shows asymptotic data collapse for the values $z_{1},z_{2}$ in Eq. (3.5).
@@
3.1 Other values of the coupling
To support our findings, we have also investigated the space-time correlator $S_{\alpha\alpha}(x,t)$ at different values of the coupling parameter $\gamma$. Only for $\gamma=1$, the point where the dynamics splits into two independent totally asymmetric processes, do we retrieve the well-known KPZ scaling of a stochastic scalar Burgers equation $z_{1}=z_{2}=3/2$, with $f_{\alpha}=f_{\rm KPZ}$ the celebrated KPZ scaling function obtained in Ref. [2]. For all other value of the coupling parameter $\gamma<1$ we find not only scaling functions distinct from $f_{\rm KPZ}$ but also different dynamical exponents, see Fig. 6. We observe a clear tendency of both $z_{1},z_{2}\rightarrow 3/2$ as $\gamma$ increases towards $\gamma=1$, the point of decoupling into separate simple exclusion processes. Also the shape of scaling functions $f_{1},f_{2}$ become closer to KPZ as $\gamma$ increases towards $1$, data not shown. The data for the regime $\gamma<\gamma_{\rm crit}$ which correspond to an isolated umbilic point ( saddle point topology) are less transparent. The critical value of the coupling $\gamma=1/4$, at which the saddle point appears, corresponds to the maximal $|z_{1}-z_{2}|$ difference. The limit $\gamma\rightarrow 0$ cannot be reached via Monte Carlo simulations since the effective MC dynamics close to $\gamma=0$ at the umbilic point becomes too slow. Further studies of this regime are warranted.
3.2 Continuum theory
To substantiate our findings we have also performed numerical simulations of a system of coupled Burgers equations [10] for the continuum fields $\phi_{\alpha}(x,t)$ with $\alpha=1,2$
$\partial_{t}\phi_{1}=\partial_{x}(2b\phi_{1}\phi_{2}+D\partial_{x}\phi_{1}+B\xi_{1}),$ (3.6)
$\partial_{t}\phi_{2}=\partial_{x}(b\phi_{1}^{2}+b\lambda\phi_{2}^{2}+D\partial_{x}\phi_{2}+B\xi_{2}),$ (3.7)
where $\xi_{\alpha}$ are uncorrelated white noise terms with unit variance. The connection with the lattice model’s continuum theory is established by the identification $\lambda=2-\gamma^{-1/2}$, see A. We consider the diagonal dynamical correlators of the continuum model
$S_{\alpha\alpha}(x,t)=\langle\phi_{\alpha}(x,t)\phi_{\alpha}(0,0)\rangle.$ (3.8)
starting from uncorrelated initial states $\langle\phi_{\alpha}(x,0),\phi_{\beta}(x^{\prime},0)\rangle=\delta_{\alpha\beta}\delta(x-x^{\prime})$ at the critical value $\lambda_{\rm crit}=0$ which corresponds to $\gamma_{\rm crit}=1/4$.
Figure 7: Time-scaling of the diagonal correlators $S_{\alpha\alpha}(0,t)$ of the system of coupled Burgers equations (3.6) and (3.7). The data is scaled along the vertical axis such that a dynamical exponent $z_{\rm KPZ}=3/2$ would correspond to asymptotically horizontal data. Dashed black lines indicate asymptotic power-laws with estimated exponents $z_{1}\approx 1.49$ and $z_{2}\approx 1.54$. Parameters: number of disretization points=$2^{20}$, $t_{\rm max}=2^{10}$, $b=3$, $\lambda=0$, $D=1$, $B=\sqrt{2}$.
We find the most robust signal of the dynamical exponent by considering the time-scaling of the correlator (3.8) at $x=0$. As shown in Fig. 7, we again find clear deviations from the KPZ dynamical exponent $z_{\rm KPZ}=3/2$ for both modes. In particular, we estimate the values of the dynamical exponents as $z_{1}\approx 1.49$ and $z_{2}\approx 1.54$. While these values for the continuum theory do not precisely match the exponent estimates for the lattice model (3.5), we note that the deviations are in the same direction, i.e. $z_{1}<z_{\rm KPZ}$ while $z_{2}>z_{\rm KPZ}$. We attribute the quantitative mismatch of the estimates to finite simulation times and different rates of convergence towards stationary values in the lattice and continuum models.
4 Conclusions
We have presented a numerical study of the dynamical structure factor for a many-body particle system with an umbilic point and two conservation laws. Both isolated and non-isolated umbilic point were considered. We have investigated the validity of the scaling hypothesis (3.4) in a basis which diagonalizes both the dynamical structure matrix $S_{\alpha\beta}(x,t)$ and the static covariance matrix. On the basis of the data we conclude that, within the hypothesis (3.4) that: (1) the dynamical exponents $z_{1},z_{2}$ and (2) the dynamical scaling functions $f_{1},f_{2}$ vary continuously with the parameter of the model $\gamma$ in a nontrivial way, i.e. the scaling functions are not merely dilated (the latter observation being in agreement with an earlier study [10]). At the umbilic point of the system of coupled Burgers equations we also find systematic deviations from the KPZ dynamical exponents for both modes. Continuously varying dynamical exponents $z_{1},z_{2}$, even though unexpected in growth processes, do not contradict the usual power counting scaling arguments (valid for one-component system), since our two-lane process is not Galilean-invariant, see [18, 20] and also A for more details.